The sun is over head at the Equator (00) on march 21st; what will be the angle of the sun at Addis Ababa (90N)



1
Expert's answer
2021-12-13T16:21:03-0500

The angle is


"\theta=\arcsin\frac{D_{ES}}{\sqrt{D_{ES}^2+(9/360·2\pi R_E)^2}},\\\space\\
\theta=\arcsin\frac{149.6·10^6}{\sqrt{(149.6·10^6)^2+(9/360·2\pi ·6378)^2}}=90°."